Determine the percentage composition of a mixture of anhydrous sodium carbonate and sodium bicarbonate from the following data
Weight of mixture weight of mixture Taken= 2 gram , loss in weight on heating= 0.124 g
 
 

Grade:11

2 Answers

Arun
5 years ago

When the mixture of sodium carbonate and sodium bicarbonate (2 g) is heated, then sodium bicarbonate decomposes to give sodium carbonate, water and carbon-di-oxide, which are evolved as gases. Therefore the loss in the weight (0.124 g) of the mixture will be due to evolution of theses gases

 

62 g (18g+44g) of gas is produced from 168 g of Sodium Bicarbonate.

Therefore, 0.124 g of gas will be produced form (0.124x168)/62 = 0.336 g of Sodium Bicarbonate.

The weight of sodium bicarbonate is 0.336 g, so, weight of sodium carbonate will be (2g-0.336g) = 1.664g.

(2).

Ashish
2 years ago
(1). When the mixture of sodium carbonate and sodium bicarbonate (2 g) is heated, then sodium bicarbonate decomposes to give sodium carbonate, water and carbon-di-oxide, which are evolved as gases. Therefore the loss in the weight (0.11 g) of the mixture will be due to evolution of theses gases as shown below:
 
 
 
62 g (18g+44g) of gas is produced from 168 g of Sodium Bicarbonate.
 
Therefore, 0.11 g of gas will be produced form (0.11x168)/62 = 0.298 g of Sodium Bicarbonate.
 
The weight of sodium bicarbonate is 0.298 g, so, weight of sodium carbonate will be (2g-0.298g) = 1.702g.
 
(2).
 
The given weight of Sodium chloride = 250 g
 
Sodium chloride is 95% pure. Therefore actual weight of Sodium chloride = (95x250)/100 = 237.5 g
 
117 g of Sodium chloride gives 142 g of Sodium sulphate.
 
Thus, 237.5 g of Sodium chloride will give (237.5x142)/117 = 288.24 g of Sodium sulphate.
 
Now, the sodium sulphate obtained will be 90% pure. Therefore, the actual weight of Sodium sulphate will be = (90x288.24)/100 = 259.46 g of sodium sulphate.
 
Was this answer helpful19
Type your question
Why do you think this post is offensive?
 It has foul language
 It is irrelevant
 Any other
SUBMIT
Ask question
Google